# Enhanced Radio Emission Between a Galaxy Cluster Pair

### v1

@@ Abstract
Interacting galaxy cluster pairs offer a unique opportunity to study the properties of the gas in the intracluster bridge connecting them. As a consequence of the encounter, both the X-ray and radio emission from the gas are expected to be enhanced by shocks and turbulence, facilitating their detection. PSZ2 G279.79+39.09 is likely an off-axis merging system at , with its two main components observed shortly after pericenter passage. In this paper, we investigate the presence of diffuse radio emission in this system. We observed the cluster pair with the MeerKAT UHF band (544–1088 MHz) for 7.5 h and uGMRT band 3 (300–500 MHz) for 8 h. These are the first targeted radio observations of this system. We discover diffuse synchrotron emission in the system, with indication of enhanced emission in the region bridging the cluster pair. The detection is based on MeerKAT UHF data, while uGMRT band 3 observations do not allow us to derive a stringent limit on the spectral index of the source. This emission is likely generated by the turbulence injected in the early stage after the cluster-cluster encounter. However, the study of its physical properties is limited by the current data. As other systems with multiple cluster components studied in recent years, this cluster pair represents an appealing target to probe nonthermal phenomena beyond the well-studied denser regions of the intracluster medium. While we report a new detection, our analysis highlights the need for multi-band observations to fully understand these sources.
show]andrea.botteon@inaf.it
1 Introduction
@@
Cluster mergers stand as the most energetic events in the Universe since the Big Bang. Upon collision, they inject turbulence and shocks into the intra-cluster medium (ICM), often giving rise to expansive synchrotron sources such as radio halos and relics (e.g. R. J. van Weeren et al., 2019). These cluster-wide nonthermal phenomena are nowadays observed in numerous massive clusters undergoing merger events (e.g. K. Knowles et al., 2022; A. Botteon et al., 2022; S. W. Duchesne et al., 2024). Halos and relics allow us to investigate the complex mechanisms at play within the ICM, where kinetic energy of large-scale motions can be channeled into electromagnetic fluctuations and particle acceleration at smaller scale (e.g. G. Brunetti & T. W. Jones, 2014). These radio sources serve also as probes for studying the astrophysical processes associated with galaxy cluster mergers and, in general, contribute to our understanding of the formation and evolution of the large-scale structure in the Universe.
@@
In recent years, significant interest has been devoted to the study of physically bound cluster pairs or systems with multiple cluster components. Observing closely spaced clusters, where individual components can be clearly identified (e.g. as distinct X-ray-emitting clumps), gives us the opportunity to study the properties of the intracluster gas, beyond the dense cores that are more easily accessible to observations (e.g. N. Werner et al., 2008; Y. Kato et al., 2015; E. Bulbul et al., 2016; H. Akamatsu et al., 2016; T. Caglar & M. Hudaverdi, 2017; T. Caglar, 2018; H. I. Kaya et al., 2019; G. E. Alvarez et al., 2022; M. S. Mirakhor et al., 2022; Y. Omiya et al., 2023; K. Migkas et al., 2025). Indeed, as a consequence of the encounter, both the X-ray and radio emissions in the bridges of matter connecting the clusters are expected to be boosted (F. Vazza et al., 2019). The detection of the radio bridges connecting the cluster pairs A399-A401 (F. Govoni et al., 2019) and A1758N-S (A. Botteon et al., 2018, 2020a) has demonstrated that relativistic particles can be accelerated by mechanisms operating on scales larger than the individual clusters, in physical regimes that are poorly explored. One hypothesis is that these radio bridges are powered by novel particle acceleration mechanisms triggered by turbulence generated in the bridges of matter connecting dynamically active clusters (G. Brunetti & F. Vazza, 2020). This model predicts that the radio bridge emission is volume filling and characterized by steep synchrotron spectra (i.e. , where ). This scenario seems supported by the follow-up observations of A399-A401 and A1758N-S (A. Botteon et al., 2020a; J. M. de Jong et al., 2022; C. D. Nunhokee et al., 2023; G. Pignataro et al., 2024b). It is however worth noting that reports of diffuse radio emission in clusters with two or more components are not limited to the cases of A399-A401 and A1758N-S. A number of targeted observations using various Square Kilometer Array (SKA) pathfinder and precursor instruments have revealed diffuse synchrotron sources located in-between other clusters (L. Gu et al., 2019; S. W. Duchesne et al., 2021b; M. Hoeft et al., 2021; T. Venturi et al., 2022; D. Hu et al., 2025; G. Pignataro et al., 2024c; C. Stuardi et al., 2025), while others have resulted in non-detections (A. Botteon et al., 2019a, b; M. Brüggen et al., 2021; G. Pignataro et al., 2024a). The emerging picture is that extended synchrotron sources in these environments are not ubiquitous and can exhibit a broad range of morphologies, sometimes appearing as bridges of radio emission connecting X-ray clumps, and in other cases as filamentary and/or elongated structures. Whether all these sources can be attributed to the same dynamical processes and acceleration mechanisms remains unclear, motivating further theoretical and observational work on systems with multiple cluster components.
@@
PSZ2 G279.79+39.09 (hereafter G279+39) was identified in the second Planck catalog of Sunyaev-Zel’dovich (SZ) sources Planck Collaboration XXII (2016). It forms a massive ( ) binary cluster at , in which its east and west components, identified by their X-ray emission, are separated by a projected distance of 1.3 Mpc. This system has been overlook in the literature, lacking a detailed study in any band. We recently analyzed the XMM-Newton observation available on this system and characterized the properties of the two subclusters and of the thermal emission between them (S. Döner et al., 2025). Our analysis suggests that G279+39 is a massive cluster pair undergoing an off-axis merger, observed shortly after pericenter passage, with signatures of shock-heated gas in the region connecting the two subclusters. In this paper, we present new MeerKAT UHF (J. L. Jonas, 2009) and upgraded Giant Metrewave Radio Telescope (uGMRT; Y. Gupta et al., 2017) band 3 observations of G279+39 with the goal of investigating the presence of diffuse radio emission in the system. In particular, in Section 2, we describe the data and processing of the radio observations. In Section 3, we present the results based on the analysis of our new radio images. In Sections 4 and 5, we discuss the findings of our work and summarize our conclusions, respectively.
@@ 2.1 MeerKAT
After downloading the data calibrated through the SARAO Science Data Processor (SDP) calibration pipeline, we used the facetselfcal.py111https://github.com/rvweeren/lofar_facet_selfcal script (R. J. van Weeren et al., 2021) to improve the calibration on the field. This script was extensively used to calibrate LOFAR observations and has been recently adapted and successfully applied also to MeerKAT data (see A. Botteon et al., 2024, 2025; M. Balboni et al., 2025; R. J. van Weeren et al., 2025). As a first step, residual radio frequency interference not flagged by the SDP pipeline was removed using AOFlagger (A. R. Offringa et al., 2010, 2012). Low S/N channels at the edges of the band were also discarded. The data were then averaged using Default PreProcessing Pipeline (DP3; G. van Diepen et al., 2018) and compressed with Dysco (A. R. Offringa, 2016). Due to the strong ionospheric errors, we performed a series of “scalarphase+scalarcomplexgain” direction-dependent (DD) calibration cycles on a large field-of-view (FoV) of using eight directions, partially correcting for the ionospheric distortions. To further improve the calibration, sources outside a reduced FoV of were removed, and a new DD calibration was performed using seven updated directions, successfully correcting for most of the remaining DD effects in the field. Given the relatively small angular extent of our target, we further limited the visibility data to a FoV of to enable faster re-imaging. On this smaller FoV, a final cycle of direction-independent (DI) calibration was carried out, effectively reducing residual artifacts around NVSS J113521201957, a bright radio galaxy belonging to G279+39.
Final imaging was done with WSClean v3.6 (A. R. Offringa et al., 2014), adopting a robust weighting of the visibilities of (D. S. Briggs, 1995) and the multiscale multifrequency deconvolution algorithm (A. R. Offringa & O. M. Smirnov, 2017). Our full-resolution image () has a noise of 4.7 Jy beam-1. To enhance the S/N of the faint diffuse emission, we also produced source-subtracted images at lower resolution (see Section 3). The central frequency of our MeerKAT images is 822 MHz and the assumed uncertainty on the flux density scale is 15% (e.g. S. P. Sikhosana et al., 2025).
@@
We observed G279+39 with the uGMRT in band 3 (300–500 MHz) for 8 h (project code: 47_058, PI: T. Caglar) on 2025 January 6. We employed the GMRT Wideband Backend (GWB; S. H. Reddy et al., 2017), spanning 200 MHz of bandwidth, which was covered with 4096 channels. The flux density calibrators 3C147 and 3C286 were observed for 12 min each at the start and at the end of the observation, respectively.
The processing was carried out with the Source Peeling and Atmospheric Modeling (spam; H. T. Intema et al., 2009) pipeline, which performs instrumental calibration (bandpass and flux density scale), data averaging, flagging, and direction-dependent calibration. The 200 MHz bandwidth was divided by spam into six frequency slices of 33.3 MHz each, which were processed independently. This is because spam was originally developed to processed GMRT Software Backend (GSB; J. Roy et al., 2010) data, which had a smaller bandwidth. Once calibrated, the six frequency slices were jointly imaged with WSClean similarly as described above for MeerKAT. For reference, our full-resolution image (), obtained with a robust weighting of the visibilities of , has a noise of 28 Jy beam-1. The central frequency of our uGMRT images is 400 MHz and the assumed uncertainty on the flux density scale is 10% (P. Chandra et al., 2004).
@@ 3 Results
Disentangling the contribution of discrete sources to that of the diffuse emission is a crucial but difficult task, and results can significantly change if this step is done in the image-plane or uv-plane (e.g. A. Botteon et al., 2025; K. Rajpurohit et al., 2025). Here, we adopt both approaches to assess the robustness of the procedure. In the top panel of Fig. 3, we show the result obtained by reimaging the MeerKAT data at lower resolution after subtracting from the visibility data the model of the discrete sources shown in Fig. 2 (top right panel). In the bottom panel of Fig. 3, we present the result of the image-plane subtraction. This was performed by masking all pixels above in the high-resolution image of Fig. 2 (top right panel), and subsequently convolving the image to match the resolution of the uv-plane-subtracted image for a direct comparison. Overall, the morphology of the radio contours is broadly consistent between the two methods, confirming the presence of faint extended emission in the system. Notably, diffuse emission is observed connecting the two clusters, spatially aligned with the X-ray emission bridge (Fig. 4). This is the only region in the field where there is significant diffuse radio emission in MeerKAT, whereas the uGMRT provides a non-detection, as shown in the large-FoV images of Fig. 7. The implications of this non-detection on the spectral index of the emission will be discussed at the end of the next Section.
@@ 4 Discussion
The newly discovered extended radio emission connects the two X-ray clumps (Fig. 4) and exhibits two surface brightness peaks: one in the east, near the peak of the eastern cluster X-ray emission (possibly indicating the presence a radio halo-like source), and another in the west, offset by 300 kpc from the center of the western cluster X-ray emission. These two peaks can be noticed also in the profiles shown in Fig. 5, which were derived by integrating the flux density within rectangular regions of 16′′ width along the source extension (see inset). Starting from each peak, we fitted the data points on the side opposite the other cluster with a decaying exponential model and mirrored the resulting profile toward the center of the cluster pair. These fits are not intended to have any physical meaning, but are rather used to visualize that the emission between the clusters declines more gradually, and that it is not simply due to the superposition of the two exponential profiles. This suggests that the radio emission connecting the two clusters has been enhanced during their interaction.
@@
Differently from the cases of A399-A401 and A1758N-S, which have been classified as systems that have not yet reached pericenter (e.g. I. Sakelliou & T. J. Ponman, 2004; L. P. David & J. C. Kempner, 2004), G279+39 is likely observed shortly after pericenter passage (S. Döner et al., 2025). While the presence of diffuse radio emission in the former systems is more surprising, as it requires the activation of mechanisms capable of reaccelerating relativistic particles and amplifying magnetic fields prior to the core passage (G. Brunetti & F. Vazza, 2020), the formation of extended radio emission in post-pericenter systems can be more easily explained in a standard turbulent reacceleration scenario. Indeed, in this case, turbulence is naturally generated in the wake of the interacting clusters (e.g. K. Roettiger et al., 1997, 1999; L. Iapichino et al., 2008), where it can dissipate part of its energy into nonthermal components, emitting synchrotron radiation. Similar phenomena have been observed in radio halos that exhibit surface brightness extensions due to the interaction with subclusters identified either in the optical or X-ray band (e.g. K.-T. Kim et al., 1989; A. Botteon et al., 2019a; G. Pignataro et al., 2024c). When the X-ray clumps of the interacting clusters are clearly separated and the radio emission lies in the region between them, the diffuse emission has been sometimes referred to as “radio bridge”. The prime example is represented by the extended emission located between the Coma cluster and the NGC4839 galaxy group (for recent analyses, see A. Bonafede et al., 2021; E. Churazov et al., 2021; M. S. Mirakhor et al., 2023). In this sense, the diffuse radio emission observed in G279+39 could also be referred to as a radio bridge. However, it is important to emphasize that this emission is likely generated in a different stage of the cluster-cluster encounter compared to the radio bridges observed in A399-A401 and A1758N-S. Another notable difference between G279+39 and these two cluster pairs is the separation of the subclusters: 1.3 Mpc for G279+39, 2 Mpc for A1758N-S, and 3 Mpc for A399-A401. Since these are sky-projected distances, they represent lower limits on the true 3D separations. Among these systems, only for A399-A401 have there been attempts to estimate the actual spatial separation between the two subclusters (H. Akamatsu et al., 2017; V. Bonjean et al., 2018; A. D. Hincks et al., 2022). In the case of G279+39, which is likely observed shortly after pericenter passage, the true physical separation between the two subclusters is not expected to be significantly larger than the projected value. Estimating the 3D separation between the eastern and western components will require dedicated follow-up observations, particularly spectroscopic optical data to assess the presence of a potential merger component along the line-of-sight.
@@
Given the small projected separation between the subclusters in G279+39 and the likely early phase post-pericenter passage, one might wonder whether the diffuse emission in this system can be decomposed into distinct components (e.g. halos and bridge), or whether it should be treated as a single, continuous structure. Or, to rephrase it as an interrogative: how far apart must two interacting clusters be, after pericenter passage, to be considered independent systems again? From an observational point of view, distinguishing the individual components is only possible if the data have sufficient resolution and sensitivity to resolve separate substructures (both morphologically and, eventually, spectrally). We believe that this is not the case for our MeerKAT data, and therefore would tend to consider the radio emission in G279+39 as a whole.
@@
For example, assuming an oblate (prolate) ellipsoidal geometry with semi-axes kpc and kpc, the volume of emission would be cm3 ( cm3). Assuming that the radio emission is uniform in the volume, its mean radio emissivity at 822 MHz is erg s-1 Hz-1 cm-3 ( erg s-1 Hz-1 cm-3). Here we adopted the flux density integrated above the level ( mJy); assuming the the level measurement ( mJy), the emissivity would be 20% lower. Extrapolating the results at 1.4 GHz, the emissivity would fall in the range erg s-1 Hz-1 cm-3, i.e. on the low-end of the emissivity distribution of radio halos (M. Murgia et al., 2024). For comparison, the emissivity estimated for the bridges in A399-A401 and A1758N-S are an order of magnitude lower (F. Govoni et al., 2019; A. Botteon et al., 2020a). The higher emissivity in G279+39 is in line with the fact that its diffuse emission is likely generated after the pericenter passage, when numerical simulations predict that the turbulent energy flux is higher (e.g. F. Vazza et al., 2017). The XRISM satellite is currently probing the velocity dispersion and bulk velocity in the ICM with X-ray microcalorimeter spectroscopy (XRISM Collaboration et al., 2025a, b, c, d, e). Although observationally very demanding, extending such measurements to intra-cluster bridges would be crucial to constrain the level of turbulence and, in turn, understand the efficiency of particle acceleration in these low-density environments. We again remind the reader that the estimates reported above rely on the following assumptions: that the source lies in the plane of the sky, it uniformly fills an ellipsoidal volume, and it has a spectral index of .

### v2

@@ Abstract
Context. Interacting pairs of galaxy clusters offer a unique opportunity to study the properties of the gas residing in the intracluster bridge connecting them. As a consequence of the encounter, both the X-ray and radio emission from the gas are expected to be enhanced by shocks and turbulence, facilitating their detection.
Aims. PSZ2 G279.79+39.09 is likely an off-axis merging system at with its two main cluster components observed at a projected distance of 1.3 Mpc. In this paper, we investigate the presence of diffuse radio emission associated with the system.
Methods. We observed this cluster pair with the MeerKAT UHF band (544–1088 MHz) for 7.5 h and with the uGMRT band 3 (300–500 MHz) for 8 h. These are the first targeted radio observations of this system.
Results. We discover diffuse synchrotron emission in the system, with indication of enhanced emission in the region bridging the cluster pair. The detection is based on the MeerKAT UHF data, while the uGMRT band 3 observation does not allow us to derive a stringent limit on the spectral index of the source. This emission is likely generated by the turbulence injected as a consequence of the cluster-cluster encounter. However, the study of its physical properties is limited by the observations currently available on the target. If the two clusters have not yet collided, this emission would resemble the radio bridges observed in A399-A401 and A1758N-S.
Conclusions. As other systems with multiple cluster components studied in recent years, the analyzed cluster pair represents an appealing target to investigate the presence of nonthermal phenomena beyond the well-studied denser regions of the intracluster medium. While in this work we presented a new detection, our analysis underlines the need for multi-band observations to fully understand these kinds of sources.
Key Words.:
galaxies: clusters: intracluster medium – galaxies: clusters: general – galaxies: clusters: individual: PSZ2 G279.79+39.09 – radiation mechanisms: nonthermal1 Introduction
@@
Cluster mergers stand as the most energetic events in the Universe since the Big Bang. Upon collision, they inject turbulence and shocks into the intracluster medium (ICM), often giving rise to expansive synchrotron sources such as radio halos and relics (e.g. van Weeren et al., 2019). These cluster-wide nonthermal phenomena are nowadays observed in numerous massive clusters undergoing merger events (e.g. Knowles et al., 2022; Botteon et al., 2022; Duchesne et al., 2024). Halos and relics allow us to investigate the complex mechanisms at play within the ICM, where kinetic energy of large-scale motions can be channeled into electromagnetic fluctuations and particle acceleration at smaller scale (e.g. Brunetti & Jones, 2014). These radio sources serve also as probes for studying the astrophysical processes associated with galaxy cluster mergers and, in general, contribute to our understanding of the formation and evolution of the large-scale structure in the Universe.
@@
In recent years, significant interest has been devoted to the study of physically bound cluster pairs or systems with multiple cluster components. Observing closely spaced clusters, where individual components can be clearly identified (e.g. as distinct X-ray-emitting clumps), gives us the opportunity to study the properties of the intracluster gas, beyond the dense cores that are more easily accessible to observations (e.g. Werner et al., 2008; Kato et al., 2015; Bulbul et al., 2016; Akamatsu et al., 2016; Caglar & Hudaverdi, 2017; Caglar, 2018; Kaya et al., 2019; Alvarez et al., 2022; Mirakhor et al., 2022; Omiya et al., 2023; Migkas et al., 2025). Indeed, as a consequence of the encounter, both the X-ray and radio emissions in the bridges of matter connecting the clusters are expected to be boosted (Vazza et al., 2019). The detection of the radio bridges connecting the cluster pairs A399-A401 (Govoni et al., 2019) and A1758N-S (Botteon et al., 2018, 2020b) has demonstrated that relativistic particles can be accelerated by mechanisms operating on scales larger than the individual clusters, in physical regimes that are poorly explored. One hypothesis is that these radio bridges are powered by novel particle acceleration mechanisms triggered by turbulence generated in the bridges of matter connecting dynamically active clusters (Brunetti & Vazza, 2020). This model predicts that the radio bridge emission is volume filling and characterized by steep synchrotron spectra (i.e. , where ). This scenario seems supported by the follow-up observations of A399-A401 and A1758N-S (Botteon et al., 2020b; de Jong et al., 2022; Nunhokee et al., 2023; Pignataro et al., 2024b). It is however worth noting that reports of diffuse radio emission in clusters with two or more components are not limited to the cases of A399-A401 and A1758N-S. A number of targeted observations using various Square Kilometer Array (SKA) pathfinder and precursor instruments have revealed diffuse synchrotron sources located in-between other clusters (Gu et al., 2019; Duchesne et al., 2021b; Hoeft et al., 2021; Venturi et al., 2022; Pignataro et al., 2024a; Hu et al., 2025; Stuardi et al., 2025), while others have resulted in non-detections (Botteon et al., 2019a, b; Brüggen et al., 2021; Pignataro et al., 2024c). The emerging picture is that extended synchrotron sources in these environments are not ubiquitous and can exhibit a broad range of morphologies, sometimes appearing as bridges of radio emission connecting X-ray clumps, and in other cases as filamentary and/or elongated structures. Whether all these sources can be attributed to the same dynamical processes and acceleration mechanisms remains unclear, motivating further theoretical and observational work on systems with multiple cluster components.
@@
PSZ2 G279.79+39.09 (hereafter G279+39) was identified in the second Planck catalog of Sunyaev-Zel’dovich (SZ) sources Planck Collaboration XXII (2016). It forms a massive ( ) binary cluster at , in which its east and west components, identified by their X-ray emission, are separated by a projected distance of 1.3 Mpc. This system has been overlooked in the literature, lacking a detailed study in any band. We recently analyzed the XMM-Newton observation available on this system and characterized the properties of the two subclusters and of the thermal emission between them (Döner et al., 2025). Our analysis suggests that G279+39 is a massive cluster pair exhibiting X-ray asymmetries, indicative of ram-pressure stripping, and high-temperatures in the region connecting the two subclusters, consistent with shock-heated or compressed gas. These features support a scenario in which the two systems are undergoing a merger with a non-zero impact parameter and are observed close to core passage (0.5 Gyr before or after). However, whether the clusters have already reached or are approaching pericenter remains uncertain, as their projected separation and observed properties are consistent with both phases seen in idealized binary merger simulations (ZuHone et al., 2018). In this paper, we present new MeerKAT UHF (Jonas, 2009) and upgraded Giant Metrewave Radio Telescope (uGMRT; Gupta et al., 2017) band 3 observations of G279+39 with the goal of investigating the presence of diffuse radio emission in the system. In particular, in Section 2, we describe the data and processing of the radio observations. In Section 3, we present the results based on the analysis of our new radio images. In Sections 4 and 5, we discuss the findings of our work and summarize our conclusions, respectively.
@@ 2.1 MeerKAT
After downloading the data calibrated through the SARAO Science Data Processor (SDP) calibration pipeline, we used the facetselfcal.py111https://github.com/rvweeren/lofar_facet_selfcal script (van Weeren et al., 2021) to improve the calibration on the field. This script was extensively used to calibrate LOFAR observations and has been recently adapted and successfully applied also to MeerKAT data (see Botteon et al., 2024, 2025; Balboni et al., 2025; van Weeren et al., 2025). As a first step, residual radio frequency interference not flagged by the SDP pipeline was removed using AOFlagger (Offringa et al., 2010, 2012). Low S/N channels at the edges of the band were also discarded. The data were then averaged using Default PreProcessing Pipeline (DP3; van Diepen et al., 2018) and compressed with Dysco (Offringa, 2016). Due to the strong ionospheric errors, we performed a series of “scalarphase+scalarcomplexgain” direction-dependent (DD) calibration cycles on a large field-of-view (FoV) of using eight directions, partially correcting for the ionospheric distortions. To further improve the calibration, sources outside a reduced FoV of were removed, and a new DD calibration was performed using seven updated directions, successfully correcting for most of the remaining DD effects in the field. Given the relatively small angular extent of our target, we further limited the visibility data to a FoV of to enable faster re-imaging. On this smaller FoV, a final cycle of direction-independent (DI) calibration was carried out, effectively reducing residual artifacts around NVSS J113521201957, a bright radio galaxy belonging to G279+39.
Final imaging was done with WSClean v3.6 (Offringa et al., 2014), adopting a robust weighting of the visibilities of (Briggs, 1995) and the multiscale multifrequency deconvolution algorithm (Offringa & Smirnov, 2017). Our full-resolution image () has a noise of 4.7 Jy beam-1. To enhance the S/N of the faint diffuse emission, we also produced source-subtracted images at lower resolution (see Section 3). The central frequency of our MeerKAT images is 822 MHz and the assumed uncertainty on the flux density scale is 15% (e.g. Sikhosana et al., 2025).
@@
We observed G279+39 with the uGMRT in band 3 (300–500 MHz) for 8 h (project code: 47_058, PI: T. Caglar) on 2025 January 6. We employed the GMRT Wideband Backend (GWB; Reddy et al., 2017), spanning 200 MHz of bandwidth, which was covered with 4096 channels. The flux density calibrators 3C147 and 3C286 were observed for 12 min each at the start and at the end of the observation, respectively.
The processing was carried out with the Source Peeling and Atmospheric Modeling (spam; Intema et al., 2009) pipeline, which performs instrumental calibration (bandpass and flux density scale), data averaging, flagging, and direction-dependent calibration. The 200 MHz bandwidth was divided by spam into six frequency slices of 33.3 MHz each, which were processed independently. This is because spam was originally developed to processed GMRT Software Backend (GSB; Roy et al., 2010) data, which had a smaller bandwidth. Once calibrated, the six frequency slices were jointly imaged with WSClean similarly as described above for MeerKAT. For reference, our full-resolution image (), obtained with a robust weighting of the visibilities of , has a noise of 28 Jy beam-1. The central frequency of our uGMRT images is 400 MHz and the assumed uncertainty on the flux density scale is 10% (Chandra et al., 2004).
@@ 3 Results
Disentangling the contribution of discrete sources to that of the diffuse emission is a crucial but difficult task, and results can significantly change if this step is done in the image-plane or uv-plane (e.g. Botteon et al., 2025; Rajpurohit et al., 2025). Here, we adopt both approaches to assess the robustness of the procedure. In the top panel of Fig. 3, we show the result obtained by reimaging the MeerKAT data at lower resolution after subtracting from the visibility data the model of the discrete sources shown in Fig. 2 (top right panel). In the bottom panel of Fig. 3, we present the result of the image-plane subtraction. This was performed by masking all pixels above in the high-resolution image of Fig. 2 (top right panel), and subsequently convolving the image to match the resolution of the uv-plane-subtracted image for a direct comparison. Overall, the morphology of the radio contours is broadly consistent between the two methods, confirming the presence of faint extended emission in the system. Notably, diffuse emission is observed connecting the two clusters, spatially aligned with the X-ray emission bridge (Fig. 4). This is the only region in the field where there is significant diffuse radio emission in MeerKAT, whereas the uGMRT provides a non-detection, as shown in the large-FoV images of Fig. 7. The implications of this non-detection on the spectral index of the emission will be discussed at the end of the next Section.
@@ 4 Discussion
The newly discovered extended radio emission connects the two X-ray clumps (Fig. 4) and exhibits two surface brightness peaks: one in the east, near the peak of the eastern cluster X-ray emission (possibly indicating the presence a radio halo-like source), and another in the west, offset by 300 kpc from the center of the western cluster X-ray emission. The origin of the latter radio peak is unclear, as it could trace a region where the radio emission is higher either due to the local dynamics of the thermal gas (which may boost the nonthermal emission), or due to the presence of relativistic plasma injected in the medium by cluster radio galaxies. The two peaks can be noticed also in the profiles shown in Fig. 5, which were derived by integrating the flux density within rectangular regions of 16′′ width along the source extension (see inset). Starting from each peak, we fitted the data points on the side opposite the other cluster with a decaying exponential model and mirrored the resulting profile toward the center of the cluster pair. These fits are not intended to have any physical meaning, but are rather used to visualize that the emission between the clusters declines more gradually, and that it is not simply due to the superposition of the two exponential profiles. This suggests that the radio emission connecting the two clusters has been enhanced during their interaction.
The faint radio emission, which closely follows the X-ray morphology of the thermal gas, likely originates from turbulence in the region between the two interacting clusters. Since it remains uncertain whether G279+39 is observed before or after pericenter passage (Döner et al., 2025), we discuss the interpretation for its origin in both dynamical scenarios.
@@
In the pre-pericenter case, G279+39 would resemble systems such as A399-A401 and A1758N-S, which have been classified as systems that have not yet reached core passage (e.g. Sakelliou & Ponman, 2004; David & Kempner, 2004; Hincks et al., 2022; Machado et al., 2024). In these cluster pairs, the presence of radio bridges connecting the two subclusters is particularly intriguing, as it implies the activation of mechanisms capable of reaccelerating relativistic particles and amplifying magnetic fields prior to the cluster-cluster collision (Brunetti & Vazza, 2020). A notable difference between G279+39 and these two cluster pairs is the separation of the subclusters: 1.3 Mpc for G279+39, 2 Mpc for A1758N-S, and 3 Mpc for A399-A401. Since these are sky-projected distances, they represent lower limits on the true 3D separations. Among these systems, only for A399-A401 have there been attempts to estimate the actual spatial separation between the two subclusters (Akamatsu et al., 2017; Bonjean et al., 2018; Hincks et al., 2022). In the case of G279+39, which is closer to pericenter passage, the true physical separation between the two subclusters is not expected to be significantly larger than the projected value. Estimating the 3D separation between the eastern and western components will require dedicated follow-up observations, particularly spectroscopic optical data to assess the presence of a potential merger component along the line-of-sight.
In the post-pericenter case, the formation of extended radio emission can be more easily explained within the framework of turbulent reacceleration. Indeed, in this case, turbulence is naturally generated in the wake of the interacting clusters (e.g. Roettiger et al., 1997, 1999; Iapichino et al., 2008), where it can dissipate part of its energy into nonthermal components, producing diffuse synchrotron emission. Similar phenomena have been observed in radio halos that exhibit surface brightness extensions due to the interaction with subclusters identified either in the optical or X-ray band (e.g. Kim et al., 1989; Botteon et al., 2019a; Pignataro et al., 2024a). When the X-ray clumps of the interacting clusters are clearly separated and the radio emission lies in the region between them, the diffuse emission has been sometimes referred to as “radio bridge”. The prime example is represented by the extended emission located between the Coma cluster and the NGC4839 galaxy group (for recent analyses, see Bonafede et al., 2021; Churazov et al., 2021; Mirakhor et al., 2023). As this is also the case in G279+39, the radio emission in this system could still be referred to as a radio bridge, even in this dynamical scenario. We therefore suggest that, whenever the term “radio bridge” is used, the dynamical stage of the cluster-cluster encounter should be specified, if possible.
@@
Given the small projected separation between the subclusters in G279+39 one might wonder whether the diffuse emission in this system can be decomposed into distinct components (e.g. halos and bridge), or whether it should be treated as a single, continuous structure. Or, to rephrase it as an interrogative: how far apart must two interacting clusters be, before or after pericenter passage, to be considered independent systems? From an observational point of view, distinguishing the individual components is only possible if the data have sufficient resolution and sensitivity to resolve separate substructures (both morphologically and, eventually, spectrally). We believe that this is not the case for our MeerKAT data, and therefore would tend to consider the radio emission in G279+39 as a whole.
@@
For example, assuming an oblate (prolate) ellipsoidal geometry with semi-axes kpc and kpc, the volume of emission would be cm3 ( cm3). Assuming that the radio emission is uniform in the volume, its mean radio emissivity at 822 MHz is erg s-1 Hz-1 cm-3 ( erg s-1 Hz-1 cm-3). Here we adopted the flux density integrated above the level ( mJy); assuming the the level measurement ( mJy), the emissivity would be 20% lower. Extrapolating the results at 1.4 GHz, the emissivity would fall in the range erg s-1 Hz-1 cm-3, i.e. on the low-end of the emissivity distribution of radio halos (Murgia et al., 2024). For comparison, the emissivity estimated for the bridges in A399-A401 and A1758N-S are an order of magnitude lower (Govoni et al., 2019; Botteon et al., 2020b). Interpreting the higher emissivity in G279+39 is challenging given its uncertain merger phase. If the system is observed before pericenter passage, the high value may suggest that radio bridges can span a broad range of emissivities (similar to radio halos) and that this could, for example, depend on the separation between the two clusters. In the post-pericenter scenario, numerical simulations predict that the turbulent energy flux is higher (e.g. Vazza et al., 2017), which may in turn lead to higher synchrotron emissivities. The XRISM satellite is currently probing the velocity dispersion and bulk velocity in the ICM with X-ray microcalorimeter spectroscopy (XRISM Collaboration et al., 2025a, b, c, d, e). Although observationally very demanding, extending such measurements to intracluster bridges would be crucial to constrain the level of turbulence and, in turn, understand the efficiency of particle acceleration in these low-density environments. Similarly, following the evolution of relativistic electrons in numerical simulations through tracer particles in the region between colliding clusters, both before and after core passage, would help determining whether the properties of the radio emission in bridges differ between the two dynamical phases. We again remind the reader that the estimates reported above rely on the following assumptions: that the source lies in the plane of the sky, it uniformly fills an ellipsoidal volume, and it has a spectral index of .